# CrossMatch: Enhance Semi-Supervised Medical Image Segmentation with Perturbation Strategies and Knowledge Distillation

### v1

@@ 3.6 Performance-Friendly Implementation
The multi-encoder and decoder architecture of CrossMatch is straightforward and intuitive, and it conveniently allows for the introduction of knowledge distillation at the output stage. However, each image-level and feature-level perturbation requires multiple forward propagations, significantly increasing training costs. In response, we introduce methods for equivalent and performance-friendly implementations.
3.6.1 Equivalent Multi-encoders
For image-level perturbations, we follow the practice of most existing works by applying data augmentation operations before the input reaches the encoder stage. Specifically, we perform weak perturbation operator $\mathcal{A}^{w}$ and strong perturbation operator $\mathcal{A}^{s}$ in parallel on CPUs to ensure that the augmentation process does not occupy additional computational time resources during model iterations, efficiently utilizing hardware parallel capabilities.
3.6.2 Equivalent Multi-decoders
As shown in Fig. 2, our CrossMatch requires four feature perturbations to produce four different outputs from two decoders. For ease of explanation, let $h^{i}=x^{u}\to e^{i}$ denote the intermediate features produced by different encoders. Considering that each Mini Batch operates independently during gradient computation, suppose $h^{i}\in\mathbbm{R}^{B\times H\times W\times C}$, we can denote $h^{i}_{w,s}\in\mathbbm{R}^{2B\times H\times W\times C}$, thereby stacking each feature perturbation in the Batch dimension to achieve more efficient computation while maintaining the independence between different feature perturbation outcomes.
@@ 3.7 The pseudocode of CrossMatch
In summary, we present a self-training framework for multiple encoders and decoders based on knowledge distillation and provide a performance-friendly implementation. Algorithm 1 provides pseudocode in PyTorch style.
@@ 4.0.1 Dataset
In this study, we utilize the 2018 Left Atrium Segmentation Challenge (LA111www.cardiacatlas.org/atriaseg2018-challenge/) as a platform to evaluate the proposed CrossMatch. The challenge provides data consisting of 3D Magnetic Resonance Imaging (MRI) scans and their corresponding left atrium segmentation masks, divided into training and validation sets in an $80/20$ ratio, with an isotropic resolution of $0.625\times 0.625\times 0.625~{}mm^{3}$. Furthermore, we extend our experimental work to the Automatic Cardiac Diagnosis Challenge (ACDC222www.creatis.insa-lyon.fr/Challenge/acdc/). To ensure a fair comparison with previous works, we follow the same experimental setup when reporting the performance on the validation set.
Algorithm 1 Pseudocode of CrossMatch.
{python}
for x_w, x_s1, x_s2 in loader_u: # x_w:aug_w(x_u),x_s1:aug_s(x_w),x_s2:aug_s(x_w) # features of perturbed images h_w, h_s = e(x_w), e(x_s2) # perturbed feature pert_wn = h_w # none drop pert_ww, pert_ws = P_w(h_w), P_s(h_w) pert_sw, pert_ss = P_w(h_s), P_s(h_s) # concat in the batch dimension perts = cat( pert_wn, pert_ww, pert_ws, pert_sw, pert_ss ) # decode and split p_wn, p_ww, p_ws, p_sw, p_ss = d(perts).chunk(5) p_s = d(e(cat(x_s1, x_s2))) # hard (one-hot) pseudo mask mask_wn = pert_wn.argmax(dim=1).detach() # loss from image-level perturbation L_ip = criterion(p_s, mask_wn.repeat(2, 1, 1)) # loss from feature-level perturbation stus = cat(p_ww, p_ws, p_sw, p_ss) L_tkd = criterion(stus, mask_wn.repeat(4, 1, 1)) # loss from TA and student distillation L_dkd1 = criterion(p_sw, p_ww.argmax(dim=1)) L_dkd2 = criterion(p_ss, p_ws.argmax(dim=1)) L_dkd = (L_dkd1 + L_dkd2) / 2.0 # final unsupervised loss L_u = (L_ip + (1-eta) * L_tkd + eta * L_dkd)/2.0
@@ 4.0.2 Implementation Details
The CrossMatch is implemented based on PyTorch and use V-Net and U-Net as the baseline network for experiments on the LA and ACDC dataset, respectively. Specifically on the LA dataset, CrossMatch is optimized using the AdamW [38] optimizer for 9000 iterations, while on the ACDC dataset, it is trained using the SGD optimizer for 300 epochs. Different batch sizes are set for different datasets, with LA at 4 and ACDC at 12, ensuring an equal number of labeled and unlabeled samples per batch. For image preprocessing, images from the LA dataset are randomly cropped to $112\times 112\times 80$, and images from the ACDC dataset ware cropped to $256\times 256$. We set $\eta=0.3$ for both datasets, and set $\tau=$0.85 and $\tau=$0.95 for LA and ACDC, respectively. For performance evaluation, the LA dataset uses a sliding window strategy to achieve comprehensive segmentation of the cardiac area, while the ACDC dataset is evaluated by merging predicted slices into a 3D image. The evaluation metrics including Dice, Jaccard, 95% Hausdorff Distance (95HD) and Average Surface Distance (ASD) are used in this paper. In all CrossMatch experiments, feature perturbations are set as standard dropout. The dropout rates for weak and strong perturbations are set at 25% and 75% respectively. The selection of dropout type and discussion on dropout rates are elaborated in Sec. 4.4.
Notably, to ensure the fairness of the experiments, our results are calculated using the final model weights rather than the best weights saved during training, which also demonstrates the stability of our method.
Table 1: Comparisons on the LA dataset. ”↑” and ”↓” indicate the larger and the smaller the better, respectively.
@@
Method Lab. Unlab. Dice(%)↑ Jaccard(%)↑ 95HD(voxel)↓ ASD(voxel)↓
@@
Table 2: Comparisons on the ACDC dataset. ”↑” and ”↓” indicate the larger and the smaller the better, respectively.
@@
Method Lab. Unlab. Dice(%)↑ Jaccard(%)↑ 95HD(voxel)↓ ASD(voxel)↓
@@
Figure 4: some visualization examples of several semi-supervised segmentation methods with 10% labeled data and ground truth on LA dataset.
@@ 4.1 Qualitative Comparison
The Fig. 4 presents some 3D visualization examples of all the compared methods and the corresponding ground truth on LA dataset. It can be observed that our CrossMatch outperforms other methods in terms of segmentation results. Particularly, our segmentation edges are smoother, with fewer misclassified voxels, more closely mirroring ground truth.
@@ 4.2 Quantitative Comparison
Table 1 summarizes the quantitative results and reveals that CrossMatch surpasses state-of-the-art (SOTA) techniques on LA dataset. When using 5% of the data with labels (4-label setting), although our Dice and Jaccard are close to those of SOTA methods, CrossMatch achieves significantly better results in the remaining evaluation metrics. Furthermore, significant performance improvements are realized in the scenarios with 8 and 16 labels. Especially using only 10% of the data with labels, CrossMatch exceeds the segmentation results obtained by fully supervised learning of V-Net on 100% of the data with labels, achieving a Dice of 91.33%.
Quantitative results on the ACDC dataset summarized in the Table 2 further demonstrates the effectiveness of CrossMatch. Particularly our method is more outstanding in terms of performance enhancement, where Dice is increased by 3.89% in the setting of 3-label. The experimental setups listed in both Table 1 and Table 2 are the same as those in [9], meaning all results are derived from the final iteration outcomes.
@@ 4.3 Computational Performance Analysis
Time (ms)↓ Method #Params (M)↓ #Flops (G)↓
@@
After 1000 iterations, as a straightforward fully supervised learning method, V-Net requires 22 ms per iteration, whereas these semi-supervised learning methods that require a combination of labeled and unlabeled data for training need more time, such as UA-MT [27], MC-Net [11] , MC-Net+ [12] and CAML [9] take 273 ms, 486 ms, 1269 ms and 1057 ms per iteration, respectively. In contrast, thanks to its streamlined structure and self-training pipeline, our CrossMatch only requires 210 ms per iteration, which is significantly lower than other semi-supervised segmentation methods, thus highlighting its efficient computational characteristics.

### v2

@@ 3.6 Performance-Friendly Implementation
CrossMatch’s multi-encoder/decoder setup is straightforward, but training costs rise due to multiple forward propagations for perturbations. Therefore, we propose effective and equivalent implementation methods.
For image-level perturbations, we apply data augmentations (weak $\mathcal{A}^{w}$ and strong $\mathcal{A}^{s}$) on CPUs before the encoder stage, leveraging parallel processing to avoid additional computation during model iterations. As shown in Fig. 2, CrossMatch uses four feature perturbations from two decoders. Let $h^{i}=x^{u}\to e^{i}$ represent intermediate features from different encoders and $h^{i}_{w,s}\in\mathbbm{R}^{2B\times H\times W\times C}$ for more efficient computation, stacking features in the Batch dimension to preserve the independence of perturbation outcomes.
@@ 3.7 The pseudocode of CrossMatch
In summary, we present a self-training framework for multiple encoders and decoders based on knowledge distillation and provide a performance-friendly implementation. Algorithm 1 provides the pseudocode of CrossMatch.
@@ 4.0.1 Dataset
In this study, we utilize the 2018 Left Atrium Segmentation Challenge (LA111www.cardiacatlas.org/atriaseg2018-challenge/) as a platform to evaluate the proposed CrossMatch. The challenge provides data consisting of 3D MRI scans and their corresponding left atrium segmentation masks, divided into training and validation sets in an $80/20$ ratio, with an isotropic resolution of $0.625\times 0.625\times 0.625~{}mm^{3}$. Furthermore, we extend our experimental work to the Automatic Cardiac Diagnosis Challenge (ACDC222www.creatis.insa-lyon.fr/Challenge/acdc/). To ensure a fair comparison with previous works, we follow the same experimental setup when reporting the performance on the validation set.
[…]
In the end, we use the ISIC 2018444https://challenge.isic-archive.com/data/#2018 dataset, which includes 3,694 dermoscopic RGB images. There are 2,594 images in the training set, 100 images in the validation set and 1,000 images in the test set. The original image sizes range from 540 $\times$ 722 to 4499 $\times$ 6748 pixels. We resize all images to 256 $\times$ 256 pixels and randomly shuffled the dataset for standardized processing. Depending on the experiment’s requirements, we randomly split the training set into 5%, 10%, and 20% subsets for further model training. These preprocessing steps ensure data consistency and provide standardized inputs.
All in all, we conduct experiments on four datasets, covering 2D and 3D segmentation of CT, MRI and RGB images, with a wide range of sample sizes, fully demonstrating the outstanding performance of our method.
Algorithm 1 Pseudocode of CrossMatch
1:Input: Unlabeled data $\mathcal{U}=\{x_{i}\}_{i=1}^{N}$, encoder $e$, decoder $d$, perturbation functions $\mathcal{P}^{w}$ and $\mathcal{P}^{s}$, criterion $\mathrm{H}(\cdot)$, hyperparameter $\eta$.
2:Output: Optimized $e$ and $d$.
3:for each batch $\{x_{w},x_{s1},x_{s2}\}$ in $\text{loader}_{u}$ do
4:     $\{h_{w},h_{s}\}\leftarrow e(\{x_{w},x_{s2}\})$triangleright Features of $x_{w},x_{s2}$
5:     $p^{w}_{n}\leftarrow h_{w}$triangleright None drop
6:     for $i\in\{w,s\}$ do
7:         for $j\in\{w,s\}$ do
8:              $p^{i}_{j}\leftarrow d(\mathcal{P}^{j}(h_{i}))$triangleright Apply feature perturbations to $h_{i}$ and decode.
9:         end for
10:     end for
11:     $\{p_{s1},p_{s2}\}\leftarrow d(e(\{x_{s1},x_{s2}\}))$
12:     Compute the $\mathcal{L}_{tkd}$ as shown in Eq. 10.
13:     Compute the $\mathcal{L}_{dkd}$ as shown in Eq. 11.
14:     Compute the $\mathcal{L}_{total}$ as shown in Eq. 12.
15:     Use the optimizer to update $e$ and $d$ through $\mathcal{L}_{{total}}$.
16:end for
@@ 4.0.2 Implementation Details
The CrossMatch is implemented based on PyTorch and uses V-Net and U-Net as the baseline networks for experiments on the (LA, Pancreas-CT) and (ACDC, ISIC 2018) datasets, respectively. On the LA dataset, CrossMatch use the AdamW [38] optimizer for 9000 iterations, and on the Pancreas-CT dataset, it needs training for 12000 iterations. On the ACDC and ISIC 2018 datasets, it uses the SGD optimizer for 300 and 200 epochs of optimization respectively. Different batch sizes are set for different datasets, with LA and Pancreas-CT at 4, ACDC and ISIC-2018 at 12, ensuring an equal number of labeled and unlabeled samples per batch. For image preprocessing, the LA and Pancreas-CT datasets are randomly cropped to $112\times 112\times 80$ and $96\times 96\times 96$, respectively. The ACDC dataset is cropped to $256\times 256$, and ISIC-2018 is adjusted to $256\times 256$. We set $\eta=0.2$ for the Pancreas-CT dataset and $\eta=0.3$ for the other datasets. For LA and Pancreas-CT, $\tau$ is set to 0.85, and for ACDC and ISIC-2018, $\tau$ is set to 0.95. For performance evaluation, the LA and Pancreas-CT datasets use a sliding window strategy to achieve comprehensive segmentation of the cardiac area, while the ACDC dataset is evaluated by merging predicted slices into a 3D image. For ISIC-2018, it is evaluated directly using 2D metrics. The evaluation metrics, including Dice, Jaccard, 95% Hausdorff Distance (95HD) and Average Surface Distance (ASD) are used in this paper. In all CrossMatch experiments, feature perturbations are set as a standard dropout. The dropout rates for weak and strong perturbations are set at 25% and 75%, respectively. The selection of dropout type and discussion on dropout rates are elaborated in Sec. 4.5.
Notably, to ensure the fairness of the experiments, our results are calculated using the final model weights rather than the best weights saved during training for all 3D tasks, which also demonstrates the stability of our method.
Figure 4: Visualizations of several semi-supervised segmentation methods with 10% and 20% labeled data and ground truth on Pancreas-CT dataset.
Table 1: Comparisons on the LA dataset. $\uparrow$” and $\downarrow$” indicate the larger and the smaller the better, respectively.
@@
Method Lab. Unlab. Dice(%$\uparrow$ Jaccard(%$\uparrow$ 95HD(voxel)$\downarrow$ ASD(voxel)$\downarrow$
@@
Table 2: Comparisons on the ACDC dataset. $\uparrow$” and $\downarrow$” indicate the larger and the smaller the better, respectively.
@@
Method Lab. Unlab. Dice(%$\uparrow$ Jaccard(%$\uparrow$ 95HD(voxel)$\downarrow$ ASD(voxel)$\downarrow$
@@
Table 3: Comparisons on the Pancreas-CT dataset. $\uparrow$” and $\downarrow$” indicate the larger and the smaller the better, respectively.
#Scans used Metrics
Method Lab. Unlab. Dice(%$\uparrow$ Jaccard(%$\uparrow$ 95HD(voxel)$\downarrow$ ASD(voxel)$\downarrow$
V-Net 6(10%) 0 42.56 36.71 42.61 11.68
V-Net 12(20%) 0 62.42 48.06 22.34 4.77
V-Net 62(All) 0 77.84 64.78 8.92 3.73
UA-MT [27] (MICCAI’19) 6(10%) 56(90%) 53.07 38.90 25.11 9.22
SASSNet [32] (MICCAI’20) 56.23 41.98 26.16 8.97
DTC [31] (AAAI’21) 59.54 45.61 16.53 3.12
UniMatch [8] (CVPR’23) 69.90 55.13 12.94 3.56
RCPS [39] (JBHI’24) 76.62 - 16.32 3.01
Ours 79.69 66.93 11.18 2.64
UA-MT [27] (MICCAI’19) 12(20%) 50(80%) 72.43 57.91 11.01 4.25
SASSNet [32] (MICCAI’20) 70.47 55.74 10.95 4.26
DTC [31] (AAAI’21) 75.94 62.41 8.25 2.21
UniMatch [8] (CVPR’23) 79.52 66.64 13.05 3.02
RCPS [39] (JBHI’24) 81.59 - 7.50 2.03
Ours 83.13 71.46 5.20 1.88
Table 4: Comparisons on the ISIC 2018 dataset. $\uparrow$” and $\downarrow$” indicate the larger and the smaller the better, respectively.
#Scans used Metrics
Method Lab. Unlab. Dice(%$\uparrow$ Jaccard(%$\uparrow$ 95HD(voxel)$\downarrow$ ASD(voxel)$\downarrow$
U-Net 129(5%) 0 60.75 49.41 82.28 38.69
U-Net 259(10%) 0 76.16 66.55 43.76 17.21
U-Net 518(20%) 0 81.18 71.54 35.73 14.22
U-Net 2594(All) 0 86.84 78.81 24.58 9.90
UA-MT [27] (MICCAI’19) 129(5%) 2465(95%) 70.22 58.77 59.91 24.96
FixMatch [6] (NeurIPS’20) 83.15 73.54 30.40 12.02
DTC [31] (AAAI’21) 76.59 66.13 49.52 20.65
UniMatch [8] (CVPR’23) 83.44 74.15 30.37 12.33
Ours 84.10 74.69 28.60 11.28
UA-MT [27] (MICCAI’19) 259(10%) 2335(90%) 77.88 68.76 46.15 17.80
FixMatch [6] (NeurIPS’20) 83.42 73.77 27.28 10.98
DTC [31] (AAAI’21) 79.29 69.36 44.49 18.46
UniMatch [8] (CVPR’23) 84.19 75.19 25.57 10.08
Ours 84.71 75.81 25.56 9.91
UA-MT [27] (MICCAI’19) 518(20%) 2076(80%) 82.21 72.47 39.46 16.09
FixMatch [6] (NeurIPS’20) 84.01 75.67 26.98 9.99
DTC [31] (AAAI’21) 83.07 73.86 31.49 12.37
UniMatch [8] (CVPR’23) 84.76 75.60 25.44 10.24
Ours 85.43 76.68 23.77 9.34
Figure 5: some visualization examples of several semi-supervised segmentation methods with 10% labeled data and ground truth on the LA dataset.
@@ 4.1 Qualitative Comparison
Fig. 5 presents some 3D visualization examples of several compared methods and the corresponding ground truth on the LA dataset. It can be observed that our CrossMatch outperforms other methods in terms of segmentation results. Particularly, our segmentation edges are smoother, with fewer misclassified voxels, and closer to the ground truth.
In Fig. 4, we visualize the results on the Pancreas-CT dataset. CrossMatch has the smallest gap compared to the ground truth, making its segmentation results more accurate. There are fewer misclassifications and errors, and the segmentation is more continuous.
@@ 4.2 Quantitative Comparison
Table 1 summarizes the quantitative results and reveals that CrossMatch surpasses state-of-the-art (SOTA) techniques on the LA dataset. When using 5% of the data with labels (4-label setting), although our Dice and Jaccard are close to those of SOTA methods, CrossMatch achieves significantly better results in the remaining evaluation metrics. Furthermore, significant performance improvements are realized in the scenarios with 8 and 16 labels. Especially using only 10% of the data with labels, CrossMatch exceeds the segmentation results obtained by fully supervised learning of V-Net on 100% of the data with labels, achieving a Dice of 91.33%.
Quantitative results on the ACDC dataset summarized in Table 2 further demonstrate the effectiveness of CrossMatch. Particularly our method is more outstanding in terms of performance enhancement, where Dice is increased by 3.89% in the setting of 3-label. The experimental setups listed in both Table 1 and Table 2 are the same as those in [9], meaning all results are derived from the final iteration outcomes.
The quantitative results of the Pancreas-CT dataset are shown in Table 3. The results indicate that our CrossMatch outperforms the SOTA methods across all evaluation metrics in every data split. Notably, when using only 10% of labeled data, the Dice is increased by 3.07% compared to the SOTA method. When using 20% of labeled data, the 95HD is improved by 2.3%.
Table 4 shows the results on the ISIC 2018 dataset. Our CrossMatch outperforms all other methods across all data splits and metrics. Notably, with 20% labeled data, CrossMatch surpasses the distance metric performance achieved by the fully supervised method.
@@ 4.3 Computational Performance Analysis
Time (ms)$\downarrow$ Method #Params (M)$\downarrow$ #Flops (G)$\downarrow$
@@
After 1000 iterations, as a straightforward, fully supervised learning method, V-Net requires 22 ms per iteration, whereas these semi-supervised learning methods that require a combination of labeled and unlabeled data for training need more time, such as UA-MT [27], MC-Net [11], MC-Net+ [12] and CAML [9] take 273 ms, 486 ms, 1269 ms and 1057 ms per iteration, respectively. In contrast, thanks to its streamlined structure and self-training pipeline, our CrossMatch only requires 210 ms per iteration, which is significantly lower than other semi-supervised segmentation methods, thus highlighting its efficient computational characteristics.
@@ 4.4 Ablation Study
4.4.1 Ablation Study on $\mathcal{L}_{dkd}$ Components
Table 6: Ablation study on $\mathcal{L}^{w}_{dkd}$ and $\mathcal{L}^{s}_{dkd}$. Comparison on the LA dataset.
Components #Labeled Metrics
$\mathcal{L}^{w}_{dkd}$ $\mathcal{L}^{s}_{dkd}$Scans used Dice(%$\uparrow$ Jaccard(%$\uparrow$ 95HD(voxel)$\downarrow$ ASD(voxel)$\downarrow$
89.26 80.71 7.35 2.17
89.28 80.72 7.45 2.15
89.80 81.58 5.94 1.77
4(5%) 89.98 81.84 5.90 1.85
86.18 76.28 10.35 2.02
85.66 76.51 9.32 1.94
89.82 81.64 6.34 1.77
8(10%) 91.33 84.11 5.29 1.53
90.98 83.53 5.78 1.85
91.46 84.33 5.50 1.63
91.25 83.98 6.39 2.07
16(20%) 91.61 84.57 5.36 1.57
Tab. 6 presents an ablation study to evaluate the impact of the loss components $\mathcal{L}^{w}_{dkd}$ and $\mathcal{L}^{s}_{dkd}$ on the performance of our model using the LA dataset. We choose the optimal parameters, and the detailed process can be found in the discussion section. To ensure the results are universal, we conduct experiments across all data splits.
When using 5%, 10% and 20% of labeled data, the addition of $\mathcal{L}^{w}_{dkd}$ and $\mathcal{L}^{s}_{dkd}$ improves Dice performance by 0.71%, 5.14%, and 0.63%, respectively. Particularly in the cases of 5% and 10%, the introduction of $\mathcal{L}^{s}_{dkd}$ significantly enhances the model’s distance metric performance due to the regularization effect of the large amount of unlabeled data. From the experimental results, it’s evident that with the combined effects of $\mathcal{L}^{w}_{dkd}$ and $\mathcal{L}^{s}_{dkd}$, our segmentation model achieves a dual benefit, leading to relatively precise segmentation.
4.4.2 Ablation Study on $\mathcal{L}_{tkd}$ Components
The $\mathcal{L}_{tkd}$ in Eq. 10 is actually the average of four feedforward streams under the supervision of $p^{w}_{n}$. To study the impact of losses from each feedforward stream on performance, we conduct ablation experiments. Specifically, we first remove the supervision of $p^{s}_{w}$ and $p^{w}_{s}$, as well as $p^{w}_{w}$ and $p^{s}_{s}$, to verify the necessity of the Cross operation. Then, we sequentially remove the supervision of each of the four feedforward streams, keeping only the remaining three. It’s important to note that we didn’t conduct experiments with only one supervision term remaining, as this would degrade the method, make the disturbance space too singular and revert it to the original FixMatch.
Table 7: Impact of $\mathcal{L}_{tkd}$ supervision components on performance on the LA dataset. $p^{i}_{j}$ checkmark indicates the application of the $p^{w}_{n}$ supervision loss component. 10% of labeled data are used for training.
Components Metrics
$p^{w}_{w}$ $p^{s}_{w}$ $p^{w}_{s}$ $p^{s}_{s}$ Dice(%$\uparrow$ Jaccard(%$\uparrow$ 95HD(voxel)$\downarrow$ ASD(voxel)$\downarrow$
90.24 82.36 8.43 2.10
90.32 82.45 6.59 1.81
90.94 83.44 7.59 1.93
89.04 80.53 7.48 1.94
88.55 79.65 8.15 1.93
91.03 83.59 5.57 1.62
91.33 84.11 5.29 1.53
Results of the melting research are shown in Tab. 7. The model performs best when all loss terms are used. Specifically, $p^{s}_{w}$ and $p^{w}_{s}$ can better capture edge details under moderate disturbance intensity, significantly improving distance metrics. Our Cross operation (using $p^{s}_{w},p^{w}_{s}$) is notably better than the Dual-Stream Perturbations in [8] (using $p^{w}_{w},p^{s}_{s}$). These results validate the effectiveness of our method.
4.4.3 Ablation Study on $\mathcal{L}_{ip}$ Components
Table 8: Impact of $\mathcal{L}_{ip}$ supervision components on performance on the LA dataset. $p^{s_{i}}_{n}$ checkmark indicates the application of the $p^{w}_{n}$ supervision loss component.
Components #Labeled Metrics
$p^{s_{1}}_{n}$ $p^{s_{2}}_{n}$Scans used Dice(%)↑ Jaccard(%)↑ 95HD(voxel)↓ ASD(voxel)↓
88.44 79.38 7.61 2.24
88.71 79.80 7.04 2.15
89.25 80.65 7.77 2.58
4(5%) 89.98 81.84 5.90 1.85
86.73 77.20 8.99 2.11
90.38 82.53 6.07 1.70
89.47 81.17 7.23 1.91
8(10%) 91.33 84.11 5.29 1.53
90.87 83.34 5.87 1.85
90.74 83.13 5.89 1.68
90.84 83.31 5.80 1.70
16(20%) 91.61 84.57 5.36 1.57
As shown in Fig. 2, our overall loss also includes pseudo-label supervision for two strongly perturbed feedforward streams. To explore the impact of these two supervisions on performance, we conduct an ablation study, and the results are shown in Tab. 8.
The results show that the model performs best when applying all strong perturbation feeds, indicating that incorporating multiple strong perturbation feeds is necessary.
4.5 Discussion
4.5.1 Impact of Supervised Loss Type
Table 9: The impact of $\mathcal{L}_{sup}$ loss types on performance on the LA dataset. 10% of labeled data are used for training. Mix represents the average of CE and Dice losses.
Metrics
Loss Type Dice(%)↑ Jaccard(%)↑ 95HD(voxel)↓ ASD(voxel)↓
CE 91.01 83.57 5.47 1.81
Dice 91.08 83.68 5.37 1.65
Mix 91.33 84.11 5.29 1.53
By default, the fully-supervised loss is the average of cross-entropy and Dice loss. To study the impact of the types of fully-supervised loss, we examine the cases of using only cross-entropy and only Dice loss, as shown in Tab. 9. The results indicate that the model performance is best with the mixed loss. Specifically, cross-entropy loss performs poorly when dealing with imbalanced classes; Dice loss is better at handling small targets but lacks in global optimization. Thus, the mixed loss that combines the strengths of both shows greater robustness and stability in various scenarios.
4.5.2 Impact of Decoder Distillation Loss Type
Table 10: The impact of $\mathcal{L}_{dkd}$ loss types on performance on the LA dataset.
#Labeled Metrics
Loss Type #Scans used Dice(%$\uparrow$ Jaccard(%$\uparrow$ 95HD(voxel)$\downarrow$ ASD(voxel)$\downarrow$
KL($T=1$) 89.59 81.20 6.87 2.08
KL($T=2$) 89.77 81.49 5.72 1.90
CE 88.55 79.58 7.71 1.99
Dice 4(5%) 89.98 81.84 5.90 1.85
KL($T=1$) 89.36 80.99 7.26 1.97
KL($T=2$) 88.99 80.57 7.01 1.98
CE 89.27 80.85 7.12 1.88
Dice 8(10%) 91.33 84.11 5.29 1.53
KL($T=1$) 90.74 83.13 5.99 1.70
KL($T=2$) 90.82 83.26 5.95 1.73
CE 91.03 83.62 5.78 1.84
Dice 16(20%) 91.61 84.57 5.36 1.57
In knowledge distillation, commonly used loss functions include KL divergence and cross-entropy loss. According to research by  [13], Dice loss also performs well in the field of medical image segmentation. To explore the impact of different loss functions on segmentation performance, we choose the aforementioned three for our experiments. Studies such as  [42, 18] indicate that the performance of KL divergence is affected by the temperature coefficient $T$, defined as $\sum_{j}(\mathrm{KL}(p^{w}_{j}/T\parallel p^{s}_{j}/T)/2+\mathrm{KL}(p^{s}_{j}/T\parallel p^{w}_{j}/T)/2)$, where $j\in\{w,s\}$. We set $T=1$ and $T=2$ in our experiments.
We conduct experiments on the LA dataset, with results shown in Tab. 10. The results indicate that the model performs best when using Dice loss, while KL divergence’s performance is unstable at different $T$ values, and the cross-entropy loss performs poorly. Overall, Dice loss is more suitable for medical image segmentation. Cross-entropy loss, due to its inadequate detail capturing ability, also fails to stand out in segmentation tasks.
4.5.3 Exploring Dropout Types
Fig. 7 shows the results on the type of Dropout used under different training sample ratios on the LA dataset. We investigate three different types of Dropout available in PyTorch: standard Dropout3D, AlphaDropout and FeatureAlphaDropout [43]. It is evident that using the standard Dropout3D as our feature perturbation strategy results in the best performance across all three data splits, followed by AlphaDropout and FeatureAlphaDropout. This may be due to the latter two inducing stronger feature perturbations, resulting in an increased Performance Gap between decoders, which is detrimental to the correct transfer of knowledge in knowledge distillation.
4.5.4 Effects of Hyperparameter $\eta$